# Mobility Edge for the Anderson Model on Random Regular Graphs

### 15 Mar 2026 v1

@@ 6.3. Proof of Diagonal Resolvent Convergence
We now construct the coupling between d-regular graph potentials and tree potentials used in the diagonal resolvent comparison. Sample the tree potential $(V_{u})_{u\in\mathbb{V}}$ as i.i.d. $N(0,1)$ random variables. Independently sample the graphs $(\mathcal{G}_{N})_{N\geq 1}$, each uniform over simple $d$-regular graphs on $[N]$. Let $r_{N}=\ln\ln N$, and define the event
@@
Outside the ball, meaning for indices $x\notin B_{r_{N}}$, we sample the $V^{(N)}_{x}$ variables as i.i.d. $N(0,1)$, which are additionally independent from all other variables in this construction.
@@
By Lemma 6.3 applied to $H_{N}$ (root 11) and to $H$ (root 0) with $r=r_{N}$, we obtain
@@ 7.1. Preliminaries
It is well known (see, e.g., [5, (2.2)]) that for every $z\in\mathbb{H}$, the resolvent entry $R^{\prime}_{00}(z)$ for the Anderson model on a variant of $\mathbb{T}_{d}$, where the root degree is $d$ and the degree of all other vertices is $d+1$, solves the recursive distribution equation
$\Gamma_{\eta}\stackrel{{\scriptstyle d}}{{=}}\frac{1}{V_{0}-E-\mathrm{i}\eta-t^{2}\sum_{j=1}^{d}\Gamma_{\eta,j}},$ (7.1)
where $V_{0}\sim N(0,1)$, the variables $\Gamma_{\eta,1},\dots,\Gamma_{\eta,K}$ and $\Gamma_{\eta}$ are i.i.d. and all variables on the right-hand side are independent.
@@
with $X_{\eta}$ and $Y_{\eta}$ are real. We have $Y_{\eta}>0$ almost surely and
@@
$R_{00}(E+\mathrm{i}\eta)=\frac{1}{V_{0}-E-\mathrm{i}\eta-t^{2}\sum_{j=1}^{d+1}\Gamma_{\eta}^{(j)}},$ (7.5)
where $V_{0}\sim N(0,1)$ is independent of the i.i.d. random variables $\Gamma_{\eta}^{(1)},\dots,\Gamma_{\eta}^{(d+1)}$ of $\Gamma_{\eta}$. In particular, defining
$B_{\eta}=\eta+t^{2}\sum_{j=1}^{d+1}Y_{\eta}^{(j)},$ (7.6)
@@
For all lemmas in this section, we retain the notation and hypothesis of Lemma 3.6. In particular, $E\in\mathbb{R}$ is always fixed.
@@
There exist constants $\eta_{0}(E)\in(0,1]$, $y_{*}(E)>0$, and $p_{*}(E)>0$, such that for all $\eta\in(0,\eta_{0}]$,
@@
$\mathbb{P}\bigl(\operatorname{Im}R(E+\mathrm{i}\eta)>c\bigr)\geq 2p_{0}.$ (7.9)
@@
$p_{1}\leq\mathbb{P}\left(\bigcup_{j=1}^{d+1}\{Y_{\eta}^{(j)}>y_{*}\}\right)\leq(d+1)\mathbb{P}(Y_{\eta}>y_{*}).$
@@
There exists a constant $C(t,E,d)>0$, such that for every $0<\eta\leq\eta_{0}$, the following bounds hold.
@@
has Lebesgue measure at most $2/u$. Therefore, conditioning on
$A=E+t^{2}\sum_{j=1}^{d}X_{\eta,j}\qquad\text{and}\qquad S_{\eta}=s,$
@@
$\mathbb{P}\big(|U_{\eta}|>u\big)\leq\mathbb{P}\big(|V-E|>u/2\big)+\sum_{j=1}^{d}\mathbb{P}\left(|X_{\eta,j}|>\frac{u}{2dt^{2}}\right).$
@@
$\sum_{j=1}^{d}Y_{\eta,j}>\frac{u}{2t^{2}},$
so at least one index satisfies $Y_{\eta,j}>u/(2dt^{2})$. Hence
$\mathbb{P}(S_{\eta}>u)\leq\sum_{j=1}^{d}\mathbb{P}\left(Y_{\eta,j}>\frac{u}{2dt^{2}}\right),$
@@ 7.2. Bootstrap and Conclusion
There exist constants $C_{1}(t,E,d),C_{2}(t,E,d)>0$ and $\varepsilon_{1}(t,E,d)\in(0,1]$ such that, for all $0<\varepsilon\leq\varepsilon_{1}$ and $0<\eta\leq\eta_{0}$,
$F_{\eta}(\varepsilon)\leq F_{\eta}(C_{1}\varepsilon^{1/3})^{d}+C_{2}\varepsilon^{1/3}.$ (7.16)
@@
$\mathbb{P}(S_{\eta}\leq\varepsilon^{1/3})\leq F_{\eta}(t^{-2}\varepsilon^{1/3})^{d}.$ (7.17)
@@
by (7.14), for all $\varepsilon\leq\varepsilon_{1}$ small enough that $2^{-1/2}\varepsilon^{-1/3}\geq 1$. Combining this with (7.17) and (7.18) proves (7.16) with $C_{1}=t^{-2}$. ∎
@@
There exist constants $C_{*}(t,E,d)>0$, and $\varepsilon_{*}(E,t,d)\in(0,1]$ such that, for all $0<\varepsilon\leq\varepsilon_{*}$ and $0<\eta\leq\eta_{0}$,
$\mathbb{P}(Y_{\eta}\leq\varepsilon)\leq C_{*}\varepsilon^{\beta},\qquad\beta=\frac{1}{20}.$ (7.19)
@@
Let $C_{1},C_{2},\varepsilon_{1}$ be as in Lemma 7.3. Choose
@@
so small that the sequence $(\tilde{\varepsilon}_{n})_{n\geq 0}$, defined by
$\tilde{\varepsilon}_{n+1}=\left(\frac{\tilde{\varepsilon}_{n}}{2C_{1}}\right)^{3},$ (7.20)
satisfies, for every $n\geq 0$,
$L_{n}=-\ln\tilde{\varepsilon}_{n}\geq\max\{12\ln C_{2},20\ln 2\},\qquad L_{n+1}\leq 4L_{n}.$ (7.21)
This is possible because $L_{n+1}=3L_{n}+3\ln(2C_{1})$.
@@
Since $\tilde{\varepsilon}_{0}\leq y_{*}$ and Lemma 7.1 gives $\mathbb{P}(Y_{\eta}>y_{*})\geq p_{*}$, we have
$q_{0}\leq 1-p_{*}<1,\qquad\nu_{0}>0.$ (7.22)
Moreover, because $C_{1}\tilde{\varepsilon}_{n+1}^{1/3}=\tilde{\varepsilon}_{n}/2<\tilde{\varepsilon}_{n}<\varepsilon_{1}$, Lemma 7.3 implies
$q_{n+1}\leq q_{n}^{d}+C_{2}\tilde{\varepsilon}_{n+1}^{1/3}.$ (7.23)
By (7.21),
$C_{2}\tilde{\varepsilon}_{n+1}^{1/3}=\exp\left(-\frac{1}{3}L_{n+1}+\ln C_{2}\right)\leq\exp\left(-\frac{1}{4}L_{n+1}\right).$
Therefore,
$\nu_{n+1}\geq\min\left\{d\nu_{n},\frac{1}{4}L_{n+1}\right\}-\ln 2.$ (7.24)
We claim that there exists an index $n_{*}$ such that
$d\nu_{n_{*}}\geq\frac{1}{4}L_{n_{*}+1}.$ (7.25)
If not, then (7.24) would imply, for all $n$, that
$\nu_{n+1}\geq d\nu_{n}-\ln 2.$
Setting
$w_{n}=\nu_{n}+\frac{\ln 2}{d-1},$
we would have $w_{n+1}\geq dw_{n}$, hence
$w_{n}\geq d^{n}w_{0}.$
On the other hand, the negation of (7.25) gives
$\nu_{n}<\frac{L_{n+1}}{4d}\leq\frac{4^{n}L_{1}}{4d}$
by (7.21), so $w_{n}\leq C4^{n}$. Since $d\geq 20>4$, this is impossible for large $n$. Thus there exists some $n_{*}\in\mathbb{N}$ such that (7.25) holds.
Now (7.24), (7.25), and (7.21) give
$\nu_{n_{*}+1}\geq\frac{1}{4}L_{n_{*}+1}-\ln 2\geq\frac{1}{5}L_{n_{*}+1}.$
Because $L_{n_{*}+2}\leq 4L_{n_{*}+1}$, we obtain
$\nu_{n_{*}+1}\geq\frac{1}{20}L_{n_{*}+2}.$
We now prove by induction that, for all $n\geq n_{*}+1$,
$\nu_{n}\geq\beta L_{n+1},\qquad\beta=\frac{1}{20}.$ (7.26)
The base case has just been established. If (7.26) holds at index $n$, then
$d\nu_{n}\geq d\beta L_{n+1}=\frac{d}{20}L_{n+1}\geq L_{n+1}\geq\frac{1}{4}L_{n+2},$
where the last inequality again uses (7.21). Hence (7.24) and $L_{n+2}\geq 20\ln 2$ yield
$\nu_{n+1}\geq\frac{1}{4}L_{n+2}-\ln 2\geq\beta L_{n+2}.$
This proves (7.26).
Consequently, for all $n\geq n_{*}+1$,
$q_{n}\leq\tilde{\varepsilon}_{n+1}^{\beta}.$
Now let $0<\varepsilon\leq\varepsilon_{*}=\tilde{\varepsilon}_{n_{*}+2}$ and choose $n\geq n_{*}+1$ so that
@@
$F_{\eta}(\varepsilon)\leq q_{n}\leq\tilde{\varepsilon}_{n+1}^{\beta}<\varepsilon^{\beta}.$
Thus (7.19) holds with $C_{*}=1$ and $\beta=1/20$. ∎
@@
There exists a constant $C(E,t,d)>0$ such that
@@
Let $\beta=1/20$ and $\varepsilon_{*}>0$ be given by Lemma 7.4. If $0<\varepsilon\leq t^{2}\varepsilon_{*}$ and $B_{\eta}\leq\varepsilon$, then each summand in (7.6) with $1\leq j\leq d+1$ satisfies
@@
$\mathbb{P}(B_{\eta}\leq\varepsilon)\leq\prod_{j=1}^{d+1}\mathbb{P}\left(Y_{\eta}^{(j)}\leq\frac{\varepsilon}{t^{2}}\right)\leq C\varepsilon^{\beta(d+1)}.$
Because $\beta(d+1)=(d+1)/20>1$, we conclude that
$\sup_{0<\eta\leq\eta_{0}}\int_{1}^{\infty}\mathbb{P}(B_{\eta}<s^{-1})\,ds<\infty,$
which is equivalent to (7.27). ∎
@@
Together with (7.28), this proves (3.6). ∎

### 19 Mar 2026 v2

@@ 5.3. Proof of Diagonal Resolvent Convergence
We now construct the coupling between $d$-regular graph potentials and tree potentials used in the diagonal resolvent comparison. Sample the tree potential $(V_{u})_{u\in\mathbb{V}}$ as i.i.d. $N(0,1)$ random variables. Independently sample the graphs $(\mathcal{G}_{N})_{N\geq 1}$, each uniform over simple $d$-regular graphs on $[N]$. Let $r_{N}=\lfloor\ln\ln N\rfloor$, and define the event
@@
Outside the ball, meaning for indices $x\notin B_{r_{N}}(\mathcal{G}_{N},1)$, we sample the $V^{(N)}_{x}$ variables as i.i.d. $N(0,1)$, which are additionally independent of all other variables in this construction. On the complement of $\mathcal{T}_{N}$, we sample all $V^{(N)}_{x}$ variables as i.i.d. $N(0,1)$, independently from the tree potential.
@@
By Lemma 5.3 applied to $H_{N}$ (root 11) and to $H_{d,t}$ (root 0) with $r=r_{N}$, we obtain
@@
6. Bounds for the Eigenvalue Counting Function
This section is dedicated to the proof of Lemma 3.5. The main input is the convergence of the Stieltjes transform of $W_{N}$ to the Stieltjes transform of the density of states measure of $H_{d,t}$ on the Bethe lattice. In the Gaussian setting, this follows directly from the diagonal resolvent convergence proved in Lemma 3.4, together with the $s=1$ case of Lemma 4.1. Finally, since convergence of Stieltjes transforms implies convergence of the associated spectral measures, the desired lower bound on the eigenvalue counting function follows from a lower bound on the integrated density of states of $H_{d,t}$, which we show in Appendix B, and the upper bound follows from Wegner’s estimate.
6.1. Notation
We define
$m_{N}(z)=\frac{1}{N}\operatorname{Tr}G_{N}(z),$
Next, because $H_{d,t}$ is self-adjoint, the spectral theorem implies the existence of a random projection-valued measure $\mu$ such that $H_{d,t}=\int\lambda\,d\mu(\lambda)$. The spectral measure at the root is given by
$\mu_{0}(S)=\langle\delta_{0},\mu(S)\delta_{0}\rangle$
$\nu(S)=\mathbb{E}\big[\mu_{0}(S)\big].$
For all $z\in\mathbb{H}$, the Stieltjes transform of $\nu$ is
$m_{\mathbb{T}}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,d\nu(\lambda).$
6.2. Preliminaries
We first state an upper bound on the integrated density of states of $H_{d,t}$, which is a consequence of a well-known estimate of Wegner [8, (5.27)].
Lemma 6.1.
For every $d\geq 3$ and $t>0$, there exists a constant $C(d,t)>0$ such that for all intervals $I\subset\mathbb{R}$, we have $\nu(I)\leq C|I|$.
We also require a lower bound that complements the previous Wegner estimate. It is proved in Appendix B.
Lemma 6.2.
For every $d\geq 3$, $t>0$, and $D>0$, there exists a constant $c(d,t,D)>0$ such that for all intervals $I\subset[-D,D]$, we have $\nu(I)\geq c|I|$.
We next show convergence of Stieltjes transforms.
Proposition 6.3.
Proof.
Write $z=E+\mathrm{i}\eta$, where $\eta>0$ is fixed. By exchangeability of the vertices,
$\mathbb{E}\big[m_{N}(z)\big]=\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}\big[G_{jj}(z)\big]=\mathbb{E}\big[G_{11}(z)\big].$ (6.1)
By Lemma 3.4, there exists a coupling such that $G_{11}(z)\to R_{00}(z)$ in probability. Since $|G_{11}(z)|\leq\eta^{-1}$ and $|R_{00}(z)|\leq\eta^{-1}$ deterministically, the bounded convergence theorem yields, after using (6.1),
$\lim_{N\to\infty}\mathbb{E}\big[m_{N}(z)\big]=\lim_{N\to\infty}\mathbb{E}\big[G_{11}(z)\big]=\mathbb{E}\big[R_{00}(z)\big]=m_{\mathbb{T}}(z).$ (6.2)
It remains to control the fluctuations of $m_{N}(z)$. By the $s=1$ case of Lemma 4.1 applied to the imaginary parts, and to the corresponding averages of the real parts, there exists $C(d)>0$ such that for all $u>0$ and all sufficiently large $N$,
$\mathbb{P}\Big(\big|m_{N}(z)-\mathbb{E}\big[m_{N}(z)\big]\big|\geq u\Big)\leq 2C\exp\big(-N\eta^{4}u^{2}/(4C)\big),$
which implies $m_{N}(z)-\mathbb{E}[m_{N}(z)]\to 0$ in probability. Combining this with (6.2) completes the proof. ∎
6.3. Conclusion
Proof of Lemma 3.5.
Let
$\widehat{\mu}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}}$
denote the empirical measure of $W_{N}$. By a basic property of Stieltjes transforms, the convergence of Stieltjes transforms in probability at every $z\in\mathbb{H}$ provided by Proposition 6.3 implies that $\widehat{\mu}_{N}$ converges weakly to $\nu$ in probability (see, e.g., [15, Lemma 2.1]). Since $\nu$ is absolutely continuous, by Lemma 6.1, we have that $\widehat{\mu}_{N}(I)$ converges to $\nu(I)$ in probability for every interval $I\subset\mathbb{R}$. The conclusion now follows from Lemma 6.1 and Lemma 6.2. ∎
@@ 7.1. Preliminaries
It is well known (see, e.g., [5, (2.2)]) that for every $z\in\mathbb{H}$, the resolvent entry $R^{\prime}_{00}(z)$ for the Anderson model on a variant of $\mathbb{T}_{d}$, where the root degree is d1d-1 and the degree of all other vertices is $d$, solves the recursive distribution equation
$\Gamma_{\eta}\stackrel{{\scriptstyle d}}{{=}}\frac{1}{V_{0}-E-\mathrm{i}\eta-t^{2}\sum_{j=1}^{d-1}\Gamma_{\eta,j}},$ (7.1)
where $V_{0}\sim N(0,1)$, the variables $\Gamma_{\eta,1},\dots,\Gamma_{\eta,d-1}$ and $\Gamma_{\eta}$ are i.i.d. and all variables on the right-hand side are independent.
@@
where $X_{\eta}$ and $Y_{\eta}$ are real. We have $Y_{\eta}>0$ almost surely and
@@
$R_{00}(E+\mathrm{i}\eta)=\frac{1}{V_{0}-E-\mathrm{i}\eta-t^{2}\sum_{j=1}^{d}\Gamma_{\eta}^{(j)}},$ (7.5)
where $V_{0}\sim N(0,1)$ is independent of the i.i.d. random variables $\Gamma_{\eta}^{(1)},\dots,\Gamma_{\eta}^{(d)}$, each distributed as $\Gamma_{\eta}$. In particular, defining
$B_{\eta}=\eta+t^{2}\sum_{j=1}^{d}Y_{\eta}^{(j)},$ (7.6)
@@
For all lemmas in this section, we retain the notation and hypothesis of Lemma 3.6. In particular, $E\in\mathbb{R}$, $d\geq 10$, and $t>0$ are always fixed. We also let $\rho$ denote the density of a mean zero, variance one Gaussian random variable. The constants in this section may depend on the constant in the hypothesis of Lemma 3.6, but we always omit this dependence in the notation.
@@
There exist constants $\eta_{0}(d,t,E)\in(0,1]$, $y_{*}(d,t,E)>0$, and $p_{*}(d,t,E)>0$, such that for all $\eta\in(0,\eta_{0}]$,
@@
$\mathbb{P}\bigl(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c\bigr)\geq 2p_{0}.$ (7.9)
@@
$p_{1}\leq\mathbb{P}\left(\bigcup_{j=1}^{d}\{Y_{\eta}^{(j)}>y_{*}\}\right)\leq d\mathbb{P}(Y_{\eta}>y_{*}).$
@@
There exists a constant $C(d,t,E)>0$, such that for every $0<\eta\leq\eta_{0}$, the following bounds hold.
@@
has Lebesgue measure at most $2/u$. Therefore, conditioning on $S_{\eta}$ and
$A=E+t^{2}\sum_{j=1}^{d-1}X_{\eta,j},$
@@
$\mathbb{P}\big(|U_{\eta}|>u\big)\leq\mathbb{P}\big(|V_{0}-E|>u/2\big)+\sum_{j=1}^{d-1}\mathbb{P}\left(|X_{\eta,j}|>\frac{u}{2(d-1)t^{2}}\right).$
@@
$\sum_{j=1}^{d-1}Y_{\eta,j}>\frac{u}{2t^{2}},$
so at least one index satisfies $Y_{\eta,j}>u/(2(d-1)t^{2})$. Hence
$\mathbb{P}(S_{\eta}>u)\leq\sum_{j=1}^{d-1}\mathbb{P}\left(Y_{\eta,j}>\frac{u}{2(d-1)t^{2}}\right),$
@@ 7.2. Bootstrap and Conclusion
There exist constants $C_{1}(d,t,E),C_{2}(d,t,E)>0$ and $\varepsilon_{1}(d,t,E)\in(0,1]$ such that, for all $0<\varepsilon\leq\varepsilon_{1}$ and $0<\eta\leq\eta_{0}$,
$F_{\eta}(\varepsilon)\leq F_{\eta}(C_{1}\varepsilon^{1/3})^{d-1}+C_{2}\varepsilon^{1/3}.$ (7.16)
@@
$\mathbb{P}(S_{\eta}\leq\varepsilon^{1/3})\leq F_{\eta}(t^{-2}\varepsilon^{1/3})^{d-1}.$ (7.17)
@@
by (7.14), for all $\varepsilon\leq\varepsilon_{1}$ small enough that $2^{-1/2}\varepsilon^{-1/3}\geq 1$. Set $C_{1}=\max\{1,t^{-2}\}$. Since $F_{\eta}$ is nondecreasing and $t^{-2}\leq C_{1}$, (7.17) implies
$\mathbb{P}(S_{\eta}\leq\varepsilon^{1/3})\leq F_{\eta}(C_{1}\varepsilon^{1/3})^{d-1}.$
Combining this with (7.18) and the bound on the middle event proves (7.16). ∎
@@
There exist constants $C_{*}(d,t,E)>0$ and $\varepsilon_{*}(d,t,E)\in(0,1]$ such that, for all $0<\varepsilon\leq\varepsilon_{*}$ and $0<\eta\leq\eta_{0}$,
$\mathbb{P}(Y_{\eta}\leq\varepsilon)\leq C_{*}\varepsilon^{\beta},\qquad\beta=\frac{31}{301}.$ (7.19)
@@
Let $C_{1},C_{2},\varepsilon_{1}$ be as in Lemma 7.3, and set
$\bar{q}=1-p_{*}\in(0,1).$
Define a deterministic sequence for $k\geq 0$ by
$r_{0}=\bar{q},\qquad r_{k+1}=\frac{1}{2}(r_{k}^{\,d-1}+r_{k}).$
Since $0<r_{k}<1$, we have $r_{k+1}<r_{k}$ for all $k$, and hence $r_{k}\rightarrow 0$ (since any limit point must satisfy the previous recursion). Choose $m\in\mathbb{N}$ such that
$r_{m}\leq\frac{1}{2}.$
Choose
@@
so small that, for the sequence defined by
$\tilde{\varepsilon}_{n+1}=\Bigl(\frac{\tilde{\varepsilon}_{n}}{C_{1}}\Bigr)^{3},\qquad L_{n}=-\ln\tilde{\varepsilon}_{n},$
the following hold:
$L_{0}\geq 300\ln C_{1},\qquad L_{1}\geq\max\{50\ln(2C_{2}),\,2\},$ (7.20)
and
$C_{2}\tilde{\varepsilon}_{k+1}^{1/3}\leq\frac{1}{2}\bigl(r_{k}-r_{k}^{\,d-1}\bigr)\qquad\text{for all }0\leq k<m.$ (7.21)
Since $L_{n+1}=3L_{n}+3\ln C_{1}$, the first condition implies
$L_{n+1}\leq\frac{301}{100}L_{n}\qquad\text{for all }n\geq 0.$ (7.22)
@@
Because $\tilde{\varepsilon}_{0}\leq y_{*}$, Lemma 7.1 gives
$q_{0}\leq\bar{q}=r_{0}.$
Also, since $C_{1}\tilde{\varepsilon}_{n+1}^{1/3}=\tilde{\varepsilon}_{n}$, Lemma 7.3 yields
$q_{n+1}\leq q_{n}^{\,d-1}+C_{2}\tilde{\varepsilon}_{n+1}^{1/3}.$ (7.23)
We claim that
$q_{k}\leq r_{k}\qquad\text{for all }0\leq k\leq m.$
This is true for $k=0$. If $k<m$ and $q_{k}\leq r_{k}$, then by (7.21),
$q_{k+1}\leq q_{k}^{\,d-1}+C_{2}\tilde{\varepsilon}_{k+1}^{1/3}\leq r_{k}^{\,d-1}+\frac{1}{2}\bigl(r_{k}-r_{k}^{\,d-1}\bigr)=\frac{1}{2}\bigl(r_{k}^{\,d-1}+r_{k}\bigr)=r_{k+1}.$
Thus the claim follows by induction. In particular,
$q_{m}\leq r_{m}\leq\frac{1}{2},$
and hence
$\nu_{m}\geq\ln 2.$ (7.24)
From (7.23) and $a+b\leq 2\max\{a,b\}$, and the definitions of $L_{n}$ and $\nu_{n}$, we get
$\nu_{n+1}\geq\min\Bigl\{(d-1)\nu_{n},\frac{1}{3}L_{n+1}-\ln C_{2}\Bigr\}-\ln 2.$
Since $L_{n+1}\geq L_{1}\geq 50\ln(2C_{2})$, we obtain
$\frac{1}{3}L_{n+1}-\ln C_{2}-\ln 2=\frac{1}{3}L_{n+1}-\ln(2C_{2})\geq\frac{31}{100}L_{n+1}.$
Therefore
$\nu_{n+1}\geq\min\Bigl\{(d-1)\nu_{n}-\ln 2,\frac{31}{100}L_{n+1}\Bigr\}\qquad\text{for all }n\geq 0.$ (7.25)
We claim that there exists $n_{*}\geq m$ such that
$\nu_{n_{*}}\geq\frac{31}{301}L_{n_{*}+1}.$ (7.26)
Suppose not. Then
$\nu_{n}<\frac{31}{301}L_{n+1}\qquad\text{for every }n\geq m.$
If for some $n\geq m$ we had
$(d-1)\nu_{n}-\ln 2\geq\frac{31}{100}L_{n+1},$
then (7.25) would imply
$\nu_{n+1}\geq\frac{31}{100}L_{n+1}\geq\frac{31}{301}L_{n+2},$
by (7.22), contradicting the assumption. Hence necessarily
$(d-1)\nu_{n}-\ln 2<\frac{31}{100}L_{n+1}\qquad\text{for all }n\geq m,$
and then (7.25) gives
$\nu_{n+1}\geq(d-1)\nu_{n}-\ln 2\qquad\text{for all }n\geq m.$
Iterating,
$\nu_{n}\geq(d-1)^{n-m}\Bigl(\nu_{m}-\frac{\ln 2}{d-2}\Bigr)+\frac{\ln 2}{d-2}\qquad\text{for all }n\geq m.$
By (7.24) and the assumption $d\geq 10$, the coefficient
$\nu_{m}-\frac{\ln 2}{d-2}$
is strictly positive. On the other hand, (7.22) implies
$L_{n+1}\leq\Bigl(\frac{301}{100}\Bigr)^{n-m}L_{m+1}.$
Since $d-1\geq 9>301/100$, the inequality
$(d-1)^{n-m}\Bigl(\nu_{m}-\frac{\ln 2}{d-2}\Bigr)+\frac{\ln 2}{d-2}<\frac{31}{301}\Bigl(\frac{301}{100}\Bigr)^{n-m}L_{m+1}$
cannot hold for all large $n$, contradiction. This proves (7.26).
Now (7.26) propagates. Indeed, if for some $n\geq n_{*}$,
$\nu_{n}\geq\frac{31}{301}L_{n+1},$
then, because $d-1\geq 9$ and $L_{n+1}\geq L_{1}\geq 2$,
$(d-1)\nu_{n}-\ln 2\geq\frac{279}{301}L_{n+1}-\ln 2\geq\frac{31}{100}L_{n+1}.$
Thus (7.25) implies
$\nu_{n+1}\geq\frac{31}{100}L_{n+1}\geq\frac{31}{301}L_{n+2},$
again by (7.22). By induction,
$\nu_{n}\geq\frac{31}{301}L_{n+1}\qquad\text{for all }n\geq n_{*}.$
Equivalently,
$q_{n}\leq\tilde{\varepsilon}_{n+1}^{31/301}\qquad\text{for all }n\geq n_{*}.$
Finally, let $\varepsilon_{*}=\tilde{\varepsilon}_{n_{*}}$, and fix $0<\varepsilon\leq\varepsilon_{*}$. Choose $n\geq n_{*}$ such that
@@
$F_{\eta}(\varepsilon)\leq q_{n}\leq\tilde{\varepsilon}_{n+1}^{31/301}<\varepsilon^{31/301}.$
This proves (7.19) with $C_{*}=1$. ∎
@@
There exists a constant $C(t,d,E)>0$ such that
@@
Let $\beta=31/301$ and $\varepsilon_{*}>0$ be given by Lemma 7.4. If $0<\varepsilon\leq t^{2}\varepsilon_{*}$ and $B_{\eta}\leq\varepsilon$, then each summand in (7.6) with $1\leq j\leq d$ satisfies
@@
$\mathbb{P}(B_{\eta}\leq\varepsilon)\leq\prod_{j=1}^{d}\mathbb{P}\left(Y_{\eta}^{(j)}\leq\frac{\varepsilon}{t^{2}}\right)\leq C\varepsilon^{\beta d}.$
Set $s_{0}=\max\{1,(t^{2}\varepsilon_{*})^{-1}\}$. For all $s\geq s_{0}$, we have $s^{-1}\leq t^{2}\varepsilon_{*}$, hence
$\mathbb{P}(B_{\eta}<s^{-1})\leq Cs^{-\beta d}.$
Therefore,
$\mathbb{E}[B_{\eta}^{-1}]\leq 1+\int_{1}^{\infty}\mathbb{P}(B_{\eta}<s^{-1})\,ds\leq 1+\int_{1}^{s_{0}}1\,ds+C\int_{s_{0}}^{\infty}s^{-\beta d}\,ds.$
Since $\beta d>1$, the right-hand side is finite uniformly in $0<\eta\leq\eta_{0}$. This proves (7.27). ∎
@@
Together with (7.28), this proves Lemma 3.6. ∎